### v1

# Utilize the Flow before Stepping into the Same River Twice: Certainty Represented Knowledge Flow for Refusal-Aware Instruction Tuning

Abstract
Refusal-Aware Instruction Tuning (RAIT) enables Large Language Models (LLMs) to refuse to answer unknown questions. By modifying responses of unknown questions in the training data to refusal responses such as “I don’t know”, RAIT enhances the reliability of LLMs and reduces their hallucination. Generally, RAIT modifies training samples based on the correctness of the initial LLM’s response. However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal. In this paper, we explore two primary causes of over-refusal: Static conflict emerges when the RAIT data is constructed solely on correctness criteria, causing similar samples in the LLM’s feature space to be assigned different labels (original vs. modified “I don’t know”). Dynamic conflict occurs due to the changes of LLM’s knowledge state during fine-tuning, which transforms previous unknown questions into knowns, while the training data, which is constructed based on the initial LLM, remains unchanged. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal. To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT). CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts. Second, we implement preliminary rehearsal training to characterize changes in the LLM’s knowledge state, which helps mitigate dynamic conflicts during the fine-tuning process. We conducted extensive experiments on open-ended question answering and multiple-choice question task. Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process. Source code and training data will be released at Github.
1 Introduction
Recently, Large Language Model (LLM) technology has made significant progress, becoming an important milestone towards AGI (Achiam et al. 2023; Dubey et al. 2024; Yang et al. 2024). However, current LLMs often output fabricated information, which is referred to as hallucinations (Ji et al. 2023). This phenomenon severely limits the usefulness and reliability of LLMs. An important reason for the occurrence of hallucinations is that when exposed to questions beyond their internal knowledge (i.e., unknown questions), LLMs may forcefully generate responses (Kang et al. 2024). Ideally, the reliable LLM should actively refuse to answer questions it doesn’t know to avoid incorrect responses (Wen et al. ; Li et al. 2024b). Recent studies have shown (Yang et al. 2023; Zhang et al. 2024; Xu et al. 2024; Cheng et al. 2024) that Refusal-Aware Instruction Tuning (RAIT) can enable LLMs to refuse answering questions beyond their knowledge.
Figure 1: Previous RAIT methods resulted in significant over-refusal, while our CRaFT mitigates this issue, enhancing the LLM’s reliability and helpfulness.
Generally, the RAIT process cab be described as follows: the initial LLM is prompted to answer all questions in the train set $D_{\text{src}}$. Based on the response correctness, the samples are categorized into two groups. Samples with correct responses are considered as known knowledge of the LLM. Their answers will not be modified, referred to as vanilla samples $D_{\text{van}}$. Conversely, samples with incorrect responses are considered as unknown knowledge of LLM. Their original answers are replaced with refusal responses like “I don’t know”. These samples are called IdK samples $D_{\text{idk}}$. The constructed RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$ is used to fine-tune the initial LLM, improving its ability to refuse to answer unknown questions. For simplicity, we refer to this original form of the RAIT method as Cor-RAIT.
However, (Cheng et al. 2024) reveals that the Cor-RAIT causes the fine-tuned LLM to refuse some questions that could have been answered correctly. Experiments demonstrate significant drop of LLM’s accuracy after Cor-RAIT: on the TriviaQA dataset (Joshi et al. 2017), it decreases from 45.05% to 28.57%. Similarly, on the Natural Questions dataset (Kwiatkowski et al. 2019), accuracy fell from 24.65% to 15.93%. We refer to this phenomenon as over-refusal, as shown in Figure 1.
Figure 2: Two causes of over-refusal: (a) Static conflict means the similar samples in the LLM’s feature space being assigned different labels (original vs. modified “I don’t know”). (b) Dynamic conflict arises since the LLM’s knowledge state evolves during SFT, turning initally unknown questions to knowns, while the target answer remains IdK. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal.
In addressing the over-refusal brought by Cor-RAIT, we identified two primary causes as shown in Figure 2. (1) Static Conflict: In the LLM feature space, two closely located samples might be assigned to $D_{\text{van}}$ and $D_{\text{idk}}$ under the Cor-RAIT framework. As illustrated by t-SNE in Figure 3(a), significant intersections exist between $D_{\text{van}}$ and $D_{\text{idk}}$, complicating their differentiation. These similar samples provide conflict supervision during training, impairing the LLM’s ability to distinguish between known and unknown questions, resulting in over-refusal. (2) Dynamic Conflict: This arises from overlooking the dynamic shifts in LLM’s knowledge state during training. Research (Ren et al. 2024; Ren and Sutherland 2024) shows that the knowledge state of LLMs changes during Supervised Fine-Tuning (SFT), with questions potentially shifting from unknown to known and vice versa. This phenomenon is reminiscent of Heraclitus’ saying, “no man ever steps in the same river twice.” However, current methods use static RAIT data reflecting the initial LLM’s knowledge state throughout SFT, which ignores these changes. This oversight leads to conflicts between the RAIT data and the LLM’s evolving knowledge, resulting in nefficient training and over-refusal.
To address the two problems above, we propose Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT). Our approach consists of two stages. Stage 1: Query the knowledge state and flow of LLM. First, the knowledge state of the initial LLM is probed. Compared to Cor-RAIT, we introduce response certainty on top of the LLM’s response correctness. The joint use of these two indicator variables effectively alleviates the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$. Further, to capture the dynamic changes in LLM’s knowledge state during training, we introduce a rehearsal training mechanism. This mechanism fine-tunes the initial LLM using data samples highly consistent with LLM’s internal knowledge. By not introducing any new knowledge during training(Ren et al. 2024; Kang et al. 2024), this method enables us to observe the LLM’s natural internal adjustments to its knowledge state. The difference in knowledge states between the fine-tuned and the initial LLMs helps characterize the knowledge flow within the LLM during training, aiding in the identification and resolution of dynamic conflicts. Stage 2: Refusal-aware instructions construction and tuning. Taking into account both the static knowledge state and dynamic knowledge flow, we filter out vanilla samples and IdK samples from the original RAIT data, therefore effectively reduce both of the static and dynamic conflicts. Finally, we fine-tune the initial LLM using the acquired RATI data, achieving robust overall performance.
Overall, our main contributions are as follows:
  • We conducted the in-depth analysis of static and dynamic conflicts in existing correctness-based RAIT data, revealing that they cause the trained LLMs’ mis-classification of known and unknown questions, leading to the issue of over-refusal in current RAIT methods.
  • To address these static and dynamic conflicts, we introduce CRaFT: during the construction of RAIT data, instead of solely depending on correctness, we additionally incorporate certainty to reduce static conflicts; through rehearsal training, we capture the trends of the LLM’s knowledge flow, which helps effectively mitigate dynamic conflicts. Extensive experiments show that CRaFT can effectively alleviate over-refusal and enhance the overall performance of the LLMs.
  • Given the shortcomings of existing refusal-aware metrics, we developed the Truthful Helpfulness Score (THS). This metric balances the reliability and helpfulness of LLMs, enabling a thorough and precise evaluation of their overall performance.
    Figure 3: t-SNE visualization of the LLM feature space
2 Related Work
2.1 Mitigating Hallucinations of LLMs
Researchers have developed several methods to mitigate hallucinations of LLMs, including data augmentation(Neeman et al. 2022), improving decoding strategies(Holtzman et al. 2019; Chuang et al. 2023), integrating external knowledge (Karpukhin et al. 2020), knowledge editing (Zhang, Yu, and Feng 2024; Li et al. 2024a) and honesty alignment(Zhang et al. 2024; Xu et al. 2024). Unlike the previous approach of correcting model answers, honesty alignment encourages models to say “I don’t know” when answering unknown questions. Ways to achieve honesty alignment include refusal-aware instruction tuning (RAIT) and reinforcement learning-based honesty alignment (RLHA).
2.2 Refusal-Aware Instruction Tuning
RAIT is the supervised technique that improves LLMs’ responses by training LLMs to directly respond with “I don’t know” to unknown questions. R-Tuning (Zhang et al. 2024) identifies these unknown questions by having the LLM answer each question in the training set a single time and then verifying whether the LLM’s responses match the label. In (Yang et al. 2023), the LLM is prompted to answer the same question multiple times. The target answer is then modified based on the correctness ratio of the responses. In (Wan et al. 2024), a knowledge-based verification mechanism ensures the model’s knowledge to be consistent with external trusted sources, increasing refusal accuracy and preventing the spread of false information.
2.3 Reinforcement Learning-based Honesty Alignment
RLHA improves model refusal ability by reinforcement learning with well-designed reward models. (Xu et al. 2024) proposed a novel alignment framework called RLKF, which used knowledge feedback to dynamically determine the knowledge boundary of LLM and train reliable reward models to encourage refusal responses. (Cheng et al. 2024) introduces a refusal function that learns whether the model should refuse to answer by evaluating the extent to which the input aligns with human honesty.
3 Over-Refusal: Analysis and Insights
3.1 Refusal-Aware Instruction Tuning
Given the initial LLM $\mathcal{M}_{0}$ and the instruction dataset $D_{\text{src}}$ consisting of question and answer pairs $x=(q,a)$, we select and modify data from $D_{\text{src}}$ to construct the data $D_{\text{rait}}$, composed of pairs $(q,a_{\text{rait}})$. $D_{\text{rait}}$ is then used to perform SFT on $\mathcal{M}_{0}$, resulting in the new LLM with the capability to decline unknown questions. This process is referred to as Refusal-Aware Instruction Tuning (RAIT). Existing studies (Zhang et al. 2024; Yang et al. 2023; Cheng et al. 2024) use $\mathcal{M}_{0}$ infer questions from $D_{\text{src}}$ and assess their correctness, denoted as $\mu$. As illustrated in Figure 4(a), the correctness threshold $\tau_{\mu}$ is defined firstly. Then, if $\mu<\tau_{\mu}$, the answer for the question is transformed into a new target answer, “I don’t know” and assigned to the subset $D_{\text{idk}}$, comprising what are termed as IdK samples. Conversely, if $\mu\geq\tau_{\mu}$, the original answer remains unchanged and is assigned to the subset $D_{\text{van}}$, which includes the so-called vanilla samples. The resulting RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$. This Correctness-based RAIT is referred to as Cor-RAIT.
However, LLMs exhibited significant over-refusal after Cor-RAIT, as shown in Figure 1. Subsequent sections of this chapter will analyze the causes and offer practical insights.
3.2 Static Conflicts in Cor-RAIT
During the RAIT process, LLMs learn to reject unknown samples by supervisions from $D_{\text{idk}}$. Our insight is that if the Cor-RAIT dataset contains vanilla and IdK samples that are closely positioned in the LLM’s representation space, the trained LLM may mistakenly classify similar vanilla samples (known questions) as IdK samples (unknown questions), causing over-refusals. To verify this, we analyzed the sample distributions of $D_{\text{van}}$ and $D_{\text{idk}}$. We extract latent representation of each question from the last hidden layer of the LLM. Then, t-SNE is adopted to visualize sample representations. Figure 3(a) displays the distributions of samples from the test split of MMLU dataset (Hendrycks et al. 2020) in the LLaMA-3-8B-instruct (Dubey et al. 2024) representation space, where IdK and vanilla samples have significant intersections.  [^1] We combined additional datasets and LLMs; see Appendix A.1 for similar observations.
Furthermore, we introduce the Conflict Rate for Similar Samples (CRSS) to quantitatively assess conflicts in supervision signals among similar samples in the RAIT dataset. For each sample $x_{i}$ in $D_{\text{idk}}$, we compute the cosine similarity between its question representation $r_{i}$ and the question representation $r_{j}$ of each sample $x_{j}$ in $D_{\text{van}}$. We identify and record the highest similarity value obtained. If this value exceeds the predefined similarity threshold $\tau_{\text{sim}}$, we record the occurrence. The CRSS is then calculated as:
$\text{CRSS}=\frac{\sum_{x_{i}\in D_{\text{idk}}}\mathbf{1}\left(\max_{x_{j}\in D_{\text{van}}}\text{cos}(r_{i},r_{j})>\tau_{\text{sim}}\right)}{|D_{\text{idk}}|}$
Therefore, the higher CRSS indicates more conflicting similar sample pairs, potentially leading to over-refusal. We computed the CRSS for Cor-RAIT, as shown in Figure 5. The results show that at $\tau_{\text{sim}}=0.97$, CRSS reaches significant levels across various LLM and dataset combinations, supporting earlier t-SNE findings [^2] We show more results and details in Appendix A.4..
The above analysis reveals that Cor-RAIT generates numerous similar sample pairs between $D_{\text{van}}$ and $D_{\text{idk}}$, resulting in conflicting supervision signals which leads to over-refusal. We term this “static conflict” to distinguish it from another conflict type discussed later.
Figure 4: RAIT Data Construction.
3.3 Certainty Mitigates the Static Conflicts
Theoretical analysis (see Appendix A.2) indicates a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$. This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively. To mitigate this, we propose incorporating a robust indicator variable aligned with correctness to select and construct the RAIT data. This variable should ensure that similar samples share comparable values, reducing the above misclassification. We suggest adopting the certainty of the LLM’s response as the indicator variable. Our theoretical analysis (see Appendix A.3) shows that certainty meets the above requirements. [^3] Section 4.2 discusses various methods for representing LLM response certainty. Appendix A.3 uses entropy as a measure, but our findings extend to other methods as well.
To incorporate correctness and certainty into the RAIT data selection, we developed the CorCer-RAIT framework, as shown in Figure 4(b). We visualized the sample distribution in $D_{\text{van}}$ and $D_{\text{idk}}$ using t-SNE in the LLM representation space, as shown in Figure 3(b), which shows a significant decrease in the overlap between $D_{\text{van}}$ and $D_{\text{idk}}$ compared to the Cor-RAIT in Figure 3(a). Furthermore, we calculated the CRSS for both methods, as shown in Figure 5, highlighting substantial reductions in CorCer-RAIT over Cor-RAIT. Therefore, the joint use of correctness and certainty effectively alleviates the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$.
Figure 5: CRSS of Different RAIT Samples.
3.4 Knowledge Flow and Dynamic Conflict
Research (Ren et al. 2024; Gekhman et al. 2024; Ren and Sutherland 2024) reveals that the knowledge state of LLMs evolves during the SFT process. The phenomenon, which we refer to as “knowledge flow”, can cause previously incorrectly answered questions to become correct ones during SFT. Despite this dynamic evolutions, the target answer of the training data remains static during the SFT process, which reflects the knowledge state of the initial LLM but ignores subsequent changes. We term it as “dynamic conflict” of RAIT data, which significantly contributes to the over-refusal in Cor-RAIT.
We use rehearsal training (see Chapter 4.3) to capture the LLM’s natural knowledge flow during SFT. The experiment on MMLU dataset (Hendrycks et al. 2020) and LLaMA-3-8B-instruct showed that the correctness of 69% samples, initially below 0.5, improved, thereby validating the above analysis on dynamic conflict. More experimental results are in Appendix A.5.
4 Methodology
4.1 Overview
Based on Section 3, we propose the Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT) to solve the over-refusal problem. CRaFT contains two stages, as shown in Figure 6.
Figure 6: The Framework of CRaFT.
Stage 1: Query the Knowledge State and Flow of LLM
[…]
Stage 2: Refusal-Aware Instructions Construction and Tuning
Using the knowledge state and flow from Stage 1, we select suitable samples from $D_{\text{src}}$ to construct the RAIT data $D_{\text{RAIT}}$ (see Section 4.4), which is is used to fine-tune the initial model $\mathcal{M}_{0}$.
4.2 Knowledge State Query
[…]
In the MCQA task, for a given model $\mathcal{M}$ and question $q$, the possible answers are included in $O=\{A,B,C,D\}$. We obtain the token probability of $\hat{a}$, denoted as $p(\hat{a}|q,\mathcal{M})$, where $\hat{a}\in O$. We use the probability of the target answer token to represent correctness. Certainty is calculated through negative entropy. The corresponding formulas are:
$\begin{array}[]{l}Cor(\mathcal{M},x_{i})=p(x_{i}.a|x_{i}.q,\mathcal{M}),\\ Cer(\mathcal{M},x_{i})=-\sum_{\hat{a}\in O}p(\hat{a}|x_{i}.q,\mathcal{M})\log(p(\hat{a}|x_{i}.q,\mathcal{M}))\end{array}$
In the OEQA task, following (Yang et al. 2023; Cheng et al. 2024), given a sample $x_{i}$, the LLM $\mathcal{M}$ performs inference on $x_{i}.q$ and generates responses $N$ times (with $N=10$). The generated responses $\{\hat{a}_{0},\dots,\hat{a}_{N-1}\}$ are denoted as $\hat{A}_{i}$. The generation process is carried out with a temperature of 1.0 and sampling enabled (do_sample=True).
$\begin{array}[]{l}Cor(\mathcal{M},x_{i})=\frac{1}{N}\sum_{\hat{a}_{j}\in\hat{A_{i}}}\mathbf{1}(\hat{a}_{j}=x_{i}.a)\\ Cer(\mathcal{M},x_{i})=\frac{1}{N(N-1)}\sum_{\begin{subarray}{c}\hat{a}_{j},\hat{a}_{k}\in\hat{A_{i}},j\neq k\end{subarray}}\text{cos}(E(\hat{a}_{j}),E(\hat{a}_{k}))\end{array}$
Correctness is obtained through exact match across the $N$ responses, calculating the proportion of accurate answers. Certainty is evaluated using a pretrained SentenceTransformer model  [^4] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode each response $\hat{a}_{j}$ into its vector representation $E(\hat{a}_{j})$, and the average similarity is computed between these vector embeddings (excluding diagonal elements). The correctness values range from $[0,1]$. In MCQA task, certainty ranges from [$-Entropy(N_{o})$, 0], and for OEQA, from $[0,1]$. $N_{o}$ stands for the number of options. More details about knowledge state query are in Appendix B.1.
4.3 Rehearsal Training and Knowledge Flow
[…]
Rehearsal training sample selection prioritizes those with the highest correctness and certainty. This insight is supported by (Ren et al. 2024; Kang et al. 2024; Gekhman et al. 2024), which indicates that LLMs primarily refine and activate existing knowledge rather than acquire new knowledge during SFT. We align the rehearsal training with the LLM’s internal knowledge state, ensuring a more natural and effective knowledge flow during the SFT process. More details about rehearsal training are in Appendix B.2.
4.4 Refusal-Aware Instruct Data Constuction
In contrast to Cor-RAIT, which selects RAIT samples solely based on correctness, our approach leverages four parameters $\mu$, $\sigma$, $\Delta\mu$, and $\Delta\sigma$ to characterize both the knowledge state and flow of $\mathcal{M}_{0}$. The challenge lies in making informed sample selections across these four dimensions. We propose a two-step heuristic method outlined in Algorithm 1.
Step 1
As shown in Figure 4(c), we first filter the training sample $D_{\text{src}}$ on the $\mu$ and $\Delta\mu$ plane. Setting a correctness threshold $\tau_{\mu}$, we define the vanilla candidate set $D_{\text{van}}^{1}=\{x_{i}|x_{i}.\mu\geq\tau_{\mu}\}$. For IdK candidates, unlike Cor-RAIT, we select $D_{\text{idk}}^{1}=\{x_{j}|x_{j}.\mu<\tau_{\mu}~{}and~{}x_{j}.\Delta\mu<0\}$. Samples in $D_{\text{drop}}^{1}=\{x_{k}|x_{k}.\mu<\tau_{\mu}~{}and~{}x_{k}.\Delta\mu\geq 0\}$ are discarded because their correctness is actively increasing during SFT, shifting from unknown to known, which could lead to dynamic conflicts.
Step 2
As shown in Figure 4(d), we sort both $D_{\text{van}}^{1}$ and $D_{\text{idk}}^{1}$ by certainty $\sigma$. From $D_{\text{van}}^{1}$, we select the top $N_{\text{van}}$ samples as final vanilla samples $D_{\text{van}}^{2}$, and the bottom $N_{\text{idk}}$ samples as IdK candidates of $D_{\text{idk}}^{2}$, whose answers are then modified to “I don’t know”. The samples in $D_{\text{drop}}^{2}$ are discarded. The final RAIT data $D_{\text{rait}}=D_{\text{van}}^{2}\cup D_{\text{idk}}^{2}$.
Algorithm 1 RAIT Data Construction Process
Input: $D_{\text{src}}=\{x_{0},x_{1},...,x_{N}\}$, $\tau_{\mu}$, $N_{\text{van}}$, $N_{\text{idk}}$
Output: $D_{rait}\subseteq D_{\text{src}}$
1:$D_{\text{van}}^{1}=\{x_{i}\,|\,x_{i}\in D_{\text{src}},x_{i}.\mu\geq\tau_{\mu}\}$
2:$D_{\text{idk}}^{1}=\{x_{j}\,|\,x_{j}\in D_{\text{src}},x_{j}.\mu<\tau_{\mu}\text{ and }x_{j}.\Delta\mu<0\}$
[…]
5:$D_{\text{van}}^{2}=\text{TopK}(D_{\text{van}}^{1},N_{\text{van}})$
6:$D_{\text{idk}}^{2}=\text{TopK}(D_{\text{idk}}^{1},N_{\text{idk}})$
7:for $x_{i}$ in $D_{\text{van}}^{2}$ do
8:     $x_{i}.a_{\text{rait}}$ = $x_{i}.a$
9:end for
10:for $x_{j}$ in $D_{\text{idk}}^{2}$ do
11:     $x_{j}.a_{\text{rait}}$ = “I don’t know”
12:end for
13:$D_{\text{rait}}=D_{\text{van}}^{2}\cup D_{\text{idk}}^{2}$
14:return $D_{\text{rait}}$
5 Experimental Setup
5.1 Dataset
In this paper, we evaluate two tasks: knowledge-oriented Multiple Choice Questions Answer (MCQA) and Open-ended Questions Answer (OEQA). For the MCQA task, the MMLU (Hendrycks et al. 2020) test split is used as the training set, the MMLU val split is used as the In-Domain (ID) test set, and the ARC-c (Clark et al. 2018) test split is the Out-Of-Domain (OOD) test set. For the OEQA task, the TriviaQA (Joshi et al. 2017) train split is used as the training set, the TriviaQA dev split is the ID test set, and the NQ (Kwiatkowski et al. 2019) dev split is the OOD test set. Refer to Appendix D.1 for details of the dataset.
5.2 Metric
In the post-RAIT evaluation of LLMs, each test sample is classified as correct, incorrect, or refused. We calculate the accuracy ($P_{c}$), error ($P_{w}$), and refusal rates ($P_{r}$) based on these categories. These metrics reveal trade-offs in LLM performance, prompting the essential question: How to determine the best-performing model?
Shortcomings of existing refusal-aware metrics
We conducted the in-depth analysis of existing refusal-aware metrics, identifying several design shortcomings (see Appendix C.1). We highlighted these shortcomings through constructed examples, as shown in Table 1.
Metric $\mathcal{M}_{1}$ $\mathcal{M}_{2}$ $\mathcal{M}_{3}$ $\mathcal{M}_{4}$
$P_{c}\uparrow$ 0.3 0.3 0.5 1
$P_{w}\downarrow$ 0.2 0.15 0 0
$P_{r}$ 0.5 0.55 0.5 0
$S_{\text{honesty}}$ (Yang et al. 2023) $\uparrow$ (0.8) (0.794) (1) (1)
TRUTHFUL (Cheng et al. 2024) $\uparrow$ (0.8) (0.75) 1 1
rely (Xu et al. 2024) $\uparrow$ (0.55) (0.548) 0.75 1
R-Acc (Feng et al. 2024) $\uparrow$ (0.8) (0.778) 1 1
ER (Feng et al. 2024) $\uparrow$ (0.3) (0.25) 0.5 1
A-Acc (Feng et al. 2024) $\uparrow$ (0.8) (0.75) 1 1
A-F1 (Feng et al. 2024) $\uparrow$ (0.8) (0.762) 1 1
AP (Zhang et al. 2024) $\uparrow$ (1) (1)
THS (ours) $\uparrow$ 0.1 0.15 0.5 1
Table 1: Comparison of refusal-aware metrics: The performance of constructed LLMs is $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$. However, existing metrics exhibit significant issues, as indicated by the numbers in (parentheses).
We constructed the initial model $\mathcal{M}_{0}$ along with four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$, with exhibiting exemplary properties. Performance comparisons clearly demonstrate a progressive improvement: $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$. Detailed information on $\mathcal{M}_{0}$ through $\mathcal{M}_{4}$ can be found in Appendix C.2. However, existing metrics exhibit significant issues: $S_{\text{honesty}}$ (Yang et al. 2023) inaccurately rates $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$ and equates $\mathcal{M}_{3}$ with $\mathcal{M}_{4}$. TRUTHFUL (Cheng et al. 2024) mistakenly favor $\mathcal{M}_{1}$ over $\mathcal{M}_{2}$. R-Acc, ER, A-Acc and A-F1 (Feng et al. 2024) also incorrectly ranks $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$. AP (Zhang et al. 2024) fails to differentiate between $\mathcal{M}_{3}$ and $\mathcal{M}_{4}$.
Figure 7: Truthful Helpfulness Score (THS).
Our Metric: Truthful Helpfulness Score (THS)
Due to the shortcomings of existing metrics, we propose the Truthful Helpfulness Score (THS). We first establish a Cartesian coordinate system with $P_{c}$ and $P_{w}$ as axes, where point $E_{1}$ represents the coordinates of the initial LLM, and point $E_{2}$ represents the coordinates of the refined. When $E_{2}$ falls below $OE_{1}$, a larger area of triangle $\triangle OE_{1}E_{2}$ indicates a stronger model. If $E_{2}$ is above $OE_{1}$, it suggests a decline in the model’s performance. Based on this, we define THS as the ratio of the cross product of $OE_{1}$ and $OE_{2}$ to the maximum cross product value:
$\text{THS}=(\overrightarrow{OE_{2}}\times\overrightarrow{OE_{1}})/(\overrightarrow{OA}\times\overrightarrow{OE_{1}})$
The results in Table 1 clearly demonstrate the effectiveness of THS. For a more detailed analysis of THS’s effectiveness, please refer to Appendix C.3.
5.3 Baselines
To verify the effectiveness of CRaFT, we compared it with the current mainstream methods: Init-Basic: We use the initial LLM, based on common question-answering prompts to instruct the LLM to answer questions. Init-Refuse: Compared to Init-Basic, we add instructions like “If you do not know the answer, please respond with ‘I don’t know.’” in the prompt. Van-Tuning: We randomly select $N_{\text{van}}+N_{\text{idk}}$ samples from $D_{\text{src}}$ for instruct-tuning directly, without sample modification. Cor-RAIT: We implemented this method following the settings in (Zhang et al. 2024), which is based on the correctness of the reply to filter and modify RAIT data. The prompts involved in each baseline are detailed in the Appendix D.2.
5.4 Implementation Details
In the experiments, we selected LLaMA2-7B-Chat (Touvron et al. 2023) and LLaMA3-8B-Instruct (Dubey et al. 2024) as the the initial LLM $\mathcal{M}_{0}$. In the MCQA task, we selected 5000 samples from MMLU as training data; in the OEQA tasks, we selected 10,000 samples from TriviaQA as training data. Except for Van-Tuning, where all samples are vanilla samples, in other RAIT settings, the ratio of vanilla and IdK samples is 1:4. In MCQA and OEQA tasks 5-shot and 3-shot were used in knowledge state query respectively. More details of the LLM’s knowledge state and flow query are in Appendix B. We used XTuner  [^5] https://github.com/InternLM/xtuner for all Instruct Tuning experiments, with 3 epochs and a maximum context length of 2048. In the MCQA tasks, we used LoRA, setting $r=64$, $\alpha=16$, $dropout=0.1$, and a learning rate of 2e-4; in the OEQA tasks, we used full parameter training. More details of training are in Appendix D.3. 0-shot and greedy decoding setting were used in evaluation. More details of evaluation are in Appendix D.4. OpenCompass  [^6] https://github.com/open-compass/opencompass was used for all knowledge state queries and evaluations. All experiments were conducted on NVIDIA A100-80GB GPUs.
6 Experimental Results and Analyses
6.1 Overall Performance
The experimental results on the OEQA and MCQA tasks are presented in Table 2 and 3. Under the ID setting for both types of tasks, our method outperformed four baseline models on THS, achieving the best results. Specifically, under the ID setting for OEQA, compared to the current best RAIT baseline, CRaFT improved the THS on LLaMA2-7B-Chat and LLaMA3-8B-Instruct by 3.56 and 2.72, respectively. Similarly, under the ID setting for MCQA, CRaFT improved the THS by 1.14 and 13.57, respectively. This indicates that CRaFT can significantly improve the model’s rejection capability under the ID setting. Under the OOD setting, CRaFT improved the THS on the MCQA task by 1.5 and 9.2, respectively, compared to Cor-RAIT. On the OEQA’s LLaMA2-7B-Chat, it improved by 1.76 compared to the most competitive method, Init-Refuse. Overall, CRaFT demonstrated excellent competitiveness in model generalization. Furthermore, we found that on the MCQA task, compared to other baselines, Cor-RAIT showed significant improvements under both ID and OOD settings. However, on the OEQA task, Cor-RAIT performed worse than Init-Refuse under the OOD setting. This reveals the limitations of the instruction fine-tuning method. It’s worth mentioning that Van-Tuning generally had a negative impact on the improvement of overall capability, implying that the instruction fine-tuning approach of forcing the model to answer can undermine the model’s inherent rejection capability. Therefore, although CRaFT surpassed Cor-RAIT under all tasks and settings, the improvement was limited under the OOD setting for OEQA due to training paradigm.
TriviaQA (ID) NQ (OOD)
[…]
Llama2- 7B- Chat Init-Basic 54 46 0 29.3 70.7 0
Init-Refuse 37 21.7 11.5 20.8 38.6 4.8
Van-Tuning 48.6 44.5 -3.7 18.3 50.2 -2.5
Cor-RAIT 41.3 18.3 19.7 16.2 27.6 4.7
CRaFT 38.5 12.9 23.3 15.8 22.4 6.5
Llama3- 8B- Instruct Init-Basic 66.8 33.2 0 40.3 59.7 0
Init-Refuse 53.9 20.8 12 31.1 38.6 5
Van-Tuning 55 38.1 -21.9 21 48.5 -11.7
Cor-RAIT 45.4 13.2 18.8 17.2 25.6 -0.1
CRaFT 43.5 10.9 21.5 19 27.5 0.4
Table 2: Experimental Results in OEQA Task.
MMLU (ID) ARC-c (OOD)
[…]
Llama2- 7B- Chat Init-Basic 45.6 52.8 0 53.9 46 0
Init-Refuse 36.4 38.9 3.9 44.4 35.7 2.6
Van-Tuning 47 53.1 1.2 54.5 45.5 1.2
Cor-RAIT 44.5 39.6 11.3 55.8 38.1 11.1
CRaFT 43.9 36.4 12.5 54.7 35.9 12.6
Llama3- 8B- Instruct Init-Basic 66.8 33.1 0 80.6 19.5 0
Init-Refuse 50 17 15.7 65.3 14.4 5.6
Van-Tuning 69.5 30.5 8 80.3 19.7 -1.3
Cor-RAIT 63.9 21.6 20.4 79.4 16.2 12.2
CRaFT 53.3 9.6 34 74.1 12.7 21.4
Table 3: Experimental Results in MCQA Task.
6.2 Ablation Experiments
In order to resolve the static and dynamic conflicts that lead to over-refusal, we extend Cor-RAIT to construct RAIT data using the information of correctness, certainty, and knowledge flow. We conduct sufficient ablation experiments to deeply investigate the impact of the above three factors on RAIT data selection. Due to space limitations, only the results of ablation experiments with the LLaMA2-7B-Chat model are reported, as shown in Table 4 and Table 5. Compared to Cor-RAIT, the method only introducing response certainty which named as Cor-RAIT w/o Flow achieved significant gains on the THS in the MCQA and OEQA tasks. This indicates that eliminating static conflicts can effectively mitigate the over-refusal of LLMs and this improvement is generalizable. Cor-RAIT w/o Cer only uses response correctness and knowledge flow. Similarly, experimental results show that introducing knowledge flow to filter dynamic conflicts can also maintain the factuality of the model while improving its rejection capability. Finally, CRaFT considers both static and dynamic conflicts, further enhancing performance improvement.
TriviaQA (ID) NQ (OOD)
[…]
Cor-RAIT 18.32 19.73 27.59 4.71
CRaFT w/o Flow 20.45 21.12 38.75 5.15
CRaFT w/o Cer 15.65 20.11 22.13 5.42
CRaFT 12.93 23.29 22.38 6.54
Table 4: Ablation Experiment Results in OEQA Task.
MMLU (ID) ARC-c (OOD)
[…]
Cor-RAIT 39.57 11.32 38.14 11.08
CRaFT w/o Flow 31.01 12.97 32.34 13.45
CRaFT w/o Cer 32.05 11.51 32.94 13.85
CRaFT 36.39 12.46 35.92 12.58
Table 5: Ablation Experiment Results in MCQA Task
7 Conclusion
In this paper, we reveal that the over-refusal issue in existing correctness-based RAIT methods stems from the static and dynamic conflicts within the RAIT data. To address these, we introduced CRaFT: Firstly, by incorporating response certainty, we mitigate static conflicts during the construction of RAIT data. Secondly, through rehearsal training, we capture the trends of knowledge flow within the LLM, thereby overcoming dynamic conflicts. We conducted extensive experiments on the MCQA and OEQA tasks. The results show that CRaFT significantly outperforms the existing correctness-based baselines, thereby validating the effectiveness of our method. In the future, we plan to enhance CraFT by integrating RL-based training strategies. We also plan to explore its adaptation for more complex tasks, including reasoning and multi-turn dialogue.
8 Acknowledgments
This research was supported by Shanghai Artificial Intelligence Laboratory.
Appendix A Static and Dynamic Conflicts
A.1 More t-SNE visualization of the LLM feature space
Building on our previous analysis, where we visualized the latent representations of MMLU test samples using the LLaMA3-8B-Instruct model, we extended the t-SNE analysis to include both the MMLU and TriviaQA datasets across different models, specifically LLaMA2-7B-Chat and LLaMA3-8B-Instruct. The t-SNE plots in Figure 8 highlight significant overlaps between IdK and Vanilla samples within Cor-RAIT, consistently observed across models and datasets, reinforcing our findings on representation spaces. Meanwhile, CorCer-RAIT demonstrates clearer separations, effectively reducing static conflict compared to Cor-RAIT.
Figure 8: More t-SNE visualization of the LLM feature space
A.2 Theoretical Analysis of the Relation between correctness and LLM representation
Through quantitative experimental analysis, we observed a relatively weak correlation between the latent representation and the correctness of predictions. Even small changes in the representation can lead to significant differences in correctness. We aim to delve deeper into the reasons behind this phenomenon.
We define the representation (i.e., the hidden state of the last token) as $r\in\mathbb{R}^{d}$, where $r$ is the dimension of the hidden layer. The linear layer’s parameters include a weight matrix $W\in\mathbb{R}^{V\times d}$ and a bias vector $b\in\mathbb{R}^{V}$. The logits output is given by:
$z=Wr+b$
The softmax function converts the logits into a probability distribution $p$, where the probability of the $i$-th token is:
$p_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{V}e^{z_{j}}}=\frac{e^{W_{i}r+b_{i}}}{\sum_{j=1}^{V}e^{W_{j}r+b_{j}}},\quad i=1,2,\ldots,V$
The gradient of the hidden layer representation $r$ with respect to the probability of the ground truth token $p_{t}$ is:
$\frac{\partial p_{t}}{\partial r}=p_{t}\left(W_{t}-\sum_{j=1}^{V}p_{j}W_{j}\right)$
This shows that the gradient depends on the difference between the correct token’s weight $W_{t}$ and the weighted average of all token weights, leading to larger adjustments in $r$ when this difference is significant. While the denominator in the softmax is continuous, the numerator can cause discontinuities in $p_{t}$ because small changes in $r$ can abruptly alter the predicted token, leading to sudden drops in $p_{t}$. Even small changes in $r$ can have a significant impact on $p_{t}$ due to the inherent discontinuities, often leading to highly similar samples being classified into $D_{\text{van}}$ and $D_{\text{idk}}$, respectively.
A.3 Theoretical Analysis of the Relation between Certainty and LLM Representation
Entropy serves as a measure of certainty, indicating the model’s confidence in its predictions. The entropy of the probability distribution $p$ is defined as:
$H=-\sum_{i=1}^{V}p_{i}\log p_{i}$
The gradient of entropy with respect to the hidden layer representation $r$ is:
$\frac{\partial H}{\partial r}=-\sum_{i=1}^{V}(\log p_{i}+1)p_{i}(W_{i}-p_{i}W)$
For small changes in the representation $dr$, the change in entropy can be approximated using a second-order expansion:
$dH\approx\nabla H(r)^{T}dr+\frac{1}{2}dr^{T}\nabla^{2}H(r)dr$
Since both the entropy function $H(p)$ and its gradient are continuous, small changes in representation $dr$ result in minimal changes in entropy $dH$. This aligns with our observation that minor variations in representation generally have a limited impact on certainty. Accordingly, we can ensure differences in certainty between $D_{\text{van}}$ and $D_{\text{idk}}$ to maintain distinctions in their representations, thereby mitigating static conflict.
A.4 Statistics of Static Conflicts
In addition to our previous analysis, we applied CRSS to the TriviaQA dataset, as shown in Figure 9. The results indicate that introducing certainty to some extent decreases the rate of conflicting samples in open-domain question answering. This reduction suggests that our approach effectively mitigates the negative impact of conflicting supervision signals, thereby reducing the likelihood of over-refusal.
Figure 9: CRSS of Different RAIT Samples on TriviaQA.
A.5 Statistics of Dynamic Conflicts
Dataset $n_{\text{rhs}}$ $n_{\mu<0.5}$ $n_{\Delta\mu>0}$ $n_{\mu<0.5,\Delta\mu>0}$
MMLU 200 5175 4812 3318
500 5175 5308 3252
1000 5175 4740 3571
1629 5175 4389 3551
TriviaQA 200 42457 9188 5102
1000 42457 9188 5034
3000 42457 9219 5061
8000 42457 11302 9553
Table 6: Comparison of Dataset Statistics across Different $n_{rhs}$ Values.
To capture the natural knowledge flow of the LLM during SFT, we conducted experiments using rehearsal training. We tested the LLaMA3-8B-Instruct on the MMLU dataset (14,042 samples) and the TriviaQA dataset (87,622 samples). The amount of training data for rehearsal training, denoted as $n_{\text{rhs}}$, was varied. For MMLU, $n_{\text{rhs}}$ values were set to 200, 500, 1000, and 1629, with 1629 representing the full set of training data filtered by correctness greater than 0.99. For TriviaQA, $n_{\text{rhs}}$ values were 200, 1000, 3000, and 8000. In the experiment, we denote the accuracy of the initial model on a given sample as $\mu$, the accuracy of the model after rehearsal training as $\tilde{\mu}$, and the difference between the two as $\Delta\mu$. Key metrics for the samples are summarized in Table 6. Regardless of the value of $n_{\text{rhs}}$, we found that a significant portion of samples with initial correctness less than 0.5 experienced an increase in $\mu$.
Appendix B Details of CRaFT
Query Type Instruction
Knowledge State Query
based on
TOKEN Probability
There is a single choice question about {Task}. Answer the question by replying A, B, C or D.
Question: {Question}
A. {Context_of_A}
B. {Context_of_B}
C. {Context_of_C}
D. {Context_of_D}
Answer:{Target_Token}(The Target Token takes on the values A, B, C, and D.)
Knowledge State Query
based on
Generation
There is a single choice question about {Task}. Answer the question by replying A, B, C or D.
Question: {Question}
A. {Context_of_A}
B. {Context_of_B}
C. {Context_of_C}
D. {Context_of_D}
Answer:
Table 7: Prompt templates of the two type knowledge state query. The italic {text} in curly braces represents variables that need to be replaced.
B.1 Two Types of Knowledge State Query
Our research focuses on the MCQA and OEQA tasks, which correspond to token probability-based knowledge state queries and generation-based knowledge state queries, respectively. In the MCQA task, we use an in-context learning approach with 5-shot examples to help the model understand the task. The instruction format is shown in Table 7, where Target_Token takes on the values A, B, C, and D, allowing us to calculate the corresponding token probabilities for correctness and certainty. In the OEQA task, we use a 3-shot setup. The instruction format, as shown in Table 7, involves setting the temperature parameter to 1 and enabling do_sample for $N$ iterations of inference, followed by calculations of correctness and certainty.
B.2 Rehearsal Training
Details about rehearsal training
For both of the OEQA and MCQA, the number of training data samples are 1000. We train set for rehearsal training from $D_{\text{src}}$ according to the knowledge state $\mathcal{M}_{0}$. For OEQA task, the training data $D_{\text{rehearsal}}=\{x_{i}|x_{i}.\mu\geq 0.995~{}\text{ and }~{}x_{i}.\tau\geq 0.995\}$. For MCQA task, the training data $D_{\text{rehearsal}}=\{x_{i}|x_{i}.\mu\geq 0.99\}$.
In the MCQA tasks, we used LoRA, setting $r=64$, $\alpha=16$, $dropout=0.1$; in the OEQA tasks, we used full parameter training. The training is 3 epochs, using the cosine learning rate scheduler without warming up. The maximum learning rate for MCQA task is 2e-4. The maximum learning rate on OEQA for LLaMA2-7B-Chat is 2e-5, for LLaMA3-8B-Instruct is 1e-6.
MMLU (ID) ARC-c (OOD)
[…]
Init-Basic - 66.9 33.1 0 80.6 19.5 0
Init-Refuse - 50 17 15.6 65.3 14.4 5.6
Van-Tuning - 69.5 30.5 8 80.3 19.7 -1.3
Cor-RAIT - 63.9 21.6 20.4 66.8 16.2 12.2
CRaFT 200 54.1 12.2 27.7 74.5 14.2 15.9
500 55.8 13.6 28.4 76 15 13.8
1000 53.3 9.6 34 74.1 12.7 21.4
1629 54.7 11.4 31.2 75 14.8 13.9
Table 8: Results of Reheasal Training in MCQA Task ($\mathcal{M}_{0}$ is LLaMA3-8B-Instruct).
TriviaQA (ID) NQ (OOD)
[…]
Init-Basic - 54 46 0 29.3 70.7 0
Init-Refuse - 37 21.8 11.5 20.8 38.6 4.8
Van-Tuning - 48.6 44.5 -3.7 18.3 50.2 -2.5
Cor-RAIT - 41.3 18.3 19.7 16.2 27.6 4.7
CRaFT 200 39 13.5 23.2 16.5 24.9 6.2
1000 38.5 12.9 23.3 15.8 22.4 6.5
3000 41.1 14.7 23.9 18.1 28.2 6.4
8000 40.2 14.4 23.3 17 25 6.7
Table 9: Results of Reheasal Training in OEQA Task. ($\mathcal{M}_{0}$ is LLaMA2-7B-Chat)
Ablation experiments of rehearsal training
In our work, a rehearsal training mechanism is proposed for detecting the knowledge flow generated by LLM during the SFT process. To study the impact of different levels of training on knowledge flow detection and the final experimental results of CRaFT, we adjusted the sample size of the rehearsal training set, $n_{\text{rhs}}$. Starting from $n_{\text{rhs}}=1000$, for MCQA, we set $n_{\text{rhs}}\in\{200,500,1000,1629\}$ [^7] The number of all the samples that meet the ’rehearsal training’ requirements is 1629 for MCQA task.; for OEQA, we set $n_{\text{rhs}}\in\{200,1000,3000,8000\}$. Hyperparameters other than $n_{\text{rhs}}$ remained constant. The experimental results are shown in Table 8 and Table 9.
The results show that, compared to other methods, regardless of the size of the rehearsal training dataset, CRaFT’s THS is always the highest. This indicates that rehearsal training is not sensitive to the size of the training data and has robustness. In addition, we found that with the change in $n_{rhs}$, CRaFT does not perform the same on OEQA and MCQA. On the OEQA task, CRaFT’s THS is basically stable; while on MCQA, its THS fluctuates more noticeably. This may be due to the different difficulty levels of the datasets used in the two tasks. Objectively, the OEQA task dataset we chose is simpler than the MCQA task dataset.
Appendix C More Details about Metrics
C.1 Details and Shortcomings of existing refusal-aware metrics
Previously, we mentioned that RAIT (Refusal-Aware Instruction Tuning) has two main effects on the model: it enhances refusal-aware capabilities and alters the model’s internal knowledge state, impacting both honesty and helpfulness. A good metric should be both singular and comprehensive, meaning that we should be able to evaluate the quality of the model based on this single metric. It should effectively balance honesty and helpfulness without requiring additional hyperparameters.
Unfortunately, after conducting a systematic review of the metrics used in existing refusal-aware work, we found that none of the proposed metrics meet the above criteria in RAIT. These metrics often emphasize one specific aspect of the model’s capabilities. In this section, we will conduct a detailed analysis of the existing metrics and highlight their shortcomings.
Alignment for Honesty (Yang et al. 2023) proposes a method that observes changes in a model’s response types before (t) and after (t+1) alignment for honesty, which is the most granular analysis method in current research. We will also adopt this method in our subsequent analysis of existing approaches. The paper introduces a metric, $S_{\text{honesty}}=\frac{1}{2}(S_{\text{prudence}}+(1-S_{\text{over-consv.}}))$, which seems to be a comprehensive metric. However, during our experiments, we found that Model A performed better than Model B in both correct and incorrect responses, yet its $S_{\text{honesty}}$ score was lower than Model B’s. This discrepancy arises because the metric does not fully account for changes in the model’s internal knowledge state, particularly when some samples shift from correct to wrong or from wrong to correct. We will provide specific examples to illustrate this issue in the following sections.
(Cheng et al. 2024) introduces the TRUTHFUL RATE, which is the sum of the proportion of questions correctly answered by the model and the proportion of questions that the model correctly refuses to answer. The issue with using this method in RAIT is that it relies on the initial model to classify data into known and unknown categories, and then evaluates the refusal ability of the modified model. However, after RAIT, the model’s internal state changes, meaning the classification of known and unknown data should also be adjusted accordingly. Therefore, TRUTHFUL RATE is not suitable as a metric for RAIT.
(Feng et al. 2024) uses four metrics: Reliable Accuracy (R-Acc), Effective Reliability (ER), Abstain Accuracy (A-Acc), and Abstain F1 (A-F1). This clearly violates the principle of singularity. Additionally, like (Cheng et al. 2024), it fails to account for the conversion between correct and wrong samples, making these metrics inadequate.
(Xu et al. 2024) defines precision (Prec) as $\frac{N_{c}}{N_{c}+N_{w}}$, where $N_{c}$, $N_{r}$, and $N_{w}$ represent the counts of correct, rejected, and wrong responses, respectively. Accuracy (Acc) is defined as $\frac{N_{c}}{N}$, and truthfulness (Truth) is given by $1-\frac{N_{w}}{N}$. The answer rate (Ans) is $1-\frac{N_{r}}{N}$, representing the proportion of non-rejected responses. The overall reliability (Rely) metric is then:
$\text{Rely}=\text{Ans}\cdot\text{Truth}+(1-\text{Ans})\cdot\text{Acc}.$
We further simplified Rely to $\text{Rely}=\text{Prec}+\text{Ans}-(\text{Ans})^{2}$, under the condition that $\text{Prec}+\text{Ans}<1$. When Prec is fixed, Ans increases in the interval (0, 0.5) and decreases in the interval (0.5, 1). This shows that Rely does not necessarily increase as Ans increases, which is clearly problematic. This is clearly unreasonable, and I will provide examples in the following sections to demonstrate its issues.
R-Tuning (Zhang et al. 2024) introduces the AP metric, which only considers the accuracy of willingly answered questions. This means the focus is primarily on honesty, without adequately considering helpfulness. For instance, if the model refuses to answer most questions and only responds to a few, but does so correctly, the AP value would still be high. Therefore, this metric is also unreasonable.
In summary, the existing metrics used in RAIT are inadequate because they fail to meet the criteria of being both singular and comprehensive. While some metrics focus on specific aspects like honesty or helpfulness, they often neglect the balance between the two or require multiple measurements to assess a model’s performance. Additionally, these metrics do not fully account for changes in the model’s internal knowledge state after training, which is crucial for accurately evaluating RAIT. A truly effective metric should provide a holistic evaluation of the model’s refusal-aware capabilities without introducing additional complexity or hyperparameters.
C.2 Details of $\mathcal{M}_{0}$ to $\mathcal{M}_{4}$
We adopt the method proposed in Alignment for Honesty (Yang et al. 2023) to observe the state changes of each sample before and after model fine-tuning to evaluate the model. We assume a test set $D_{\text{test}}$ with $N=100$ samples, and define an initial model $\mathcal{M}_{0}$. For the initial model, 50 samples are answered correctly, and 50 samples are answered incorrectly, with corresponding $P_{c}$ and $P_{w}$ both being 0.5. We construct four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$. The state changes of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are shown in Table 10. $\mathcal{M}_{3}$ represents the case where all previously wrong samples are changed to refuse after the model is refined, resulting in an accuracy of 0.5, a refusal rate of 0.5, and an error rate of 0. $\mathcal{M}_{4}$ is the case where all samples become correct, resulting in an accuracy of 1.
$\mathcal{M}_{0}$ $\mathcal{M}_{t}$ $\mathcal{M}_{1}$ $\mathcal{M}_{2}$
$\mathbf{N_{C}}$ $\mathbf{N_{W}}$ $\mathbf{N_{R}}$ $\mathbf{N_{C}}$ $\mathbf{N_{W}}$ $\mathbf{N_{R}}$
$\mathbf{N_{C}}$ 30 10 10 25 10 15
$\mathbf{N_{W}}$ 0 10 40 5 5 40
$\mathbf{N_{R}}$ 0 0 0 0 0 0
Table 10: State Changes for the Refined LLMs, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, from the Initial LLM $\mathcal{M}_{0}$
C.3 Effectiveness Analysis of THS
We argue for the effectiveness of our metric from two perspectives: one is through a point-by-point analysis, and the other is by evaluating the impact of RAIT on the model.
When $E_{2}$ falls at (0, 0), the THS is 0. When $E_{2}$ falls at $(P_{c1},0)$, which corresponds to the model $\mathcal{M}_{3}$, it indicates that the model refuses to answer all incorrect questions, maximizing its refusal-aware capability. When $E_{2}$ falls at (1, 0), the model achieves peak helpfulness by correctly answering all questions. Under the THS metric, the performance of these points increases sequentially.
Further analysis shows that when $P_{c2}$ is fixed, the smaller $P_{w2}$ is, the larger the area of triangle $\triangle OE_{1}E_{2}$, resulting in a higher THS. Conversely, when $P_{w2}$ is fixed, the smaller $P_{c2}$ is, the smaller the area of triangle $OE_{1}E_{2}$, and the lower the THS. This aligns well with the evaluation standards for model performance.
From another perspective, RAIT enhances refusal-aware capability, contributing to honesty, while also altering the model’s knowledge state, impacting helpfulness. If we only consider changes in helpfulness, the impact of RAIT on the model is reflected in the vector $\overrightarrow{E_{1}E_{2}}$ in Figure 10.a, which is parallel to vector $\overrightarrow{AB}$. The closer $E_{2}$ is to A, the stronger the model’s helpfulness. If we only consider changes in refusal-aware capability, the variations in $P_{c}$ and $P_{w}$ together form the vector $\overrightarrow{E_{1}E_{2}}$. If $E_{2}$ lies below $OE_{1}$, it indicates that a higher proportion of wrong samples have been converted to refusals, significantly enhancing the model’s refusal-aware capability. However, if $E_{2}$ falls on the line $OE_{1}$, even with a high refusal rate, THS will not increase. If $E_{2}$ is above $OE_{1}$, it indicates a severe over-refusal phenomenon, leading to a negative THS. Based on the above analysis, THS comprehensively measures the model’s honesty and helpfulness.
Figure 10: Analysis of THS.
Appendix D More Details about Experiments
D.1 Datasets
In this paper, we evaluate two tasks: knowledge-oriented Multiple Choice Questions Answer (MCQA) and Open-ended Questions Answer (OEQA). For the MCQA task, the MMLU (Hendrycks et al. 2020) test split is used as the training dataset. The MMLU val split and the ARC-c (Clark et al. 2018) test split are used as the In-Domain (ID) and Out-Of-Domain (OOD) test dataset, separately. The TriviaQA (Joshi et al. 2017) train split is used as the training dataset of OEQA task, the TriviaQA dev split and the NQ (Kwiatkowski et al. 2019) dev split are the In-Domain (ID) and Out-Of-Domain (OOD) test dataset. The details of the datasets are listed in Table 11.
MCQA OEQA
Train MMLU test 14,079 TriviaQA train 87,622
ID Eval MMLU val 1,540 TriviaQA dev 11,313
OOD eval ARC-c dev 1,172 NQ dev 3,610
Table 11: Details of the Datasets.
D.2 Prompts
Prompts for knowledge state query.
Prompts for knowledge state queries on MMLU and TriviaQA datasets are shown in Table 12 and Table 13.They use 5-shot and 3-shot settings respectively.
[…]
Instruction There is a single choice question about {Task}. Answer the question by replying A, B, C or D. Question: {Question} A. {Content_of_A} B. {Content_of_B} C. {Content_of_C} D. {Content_of_D} Answer:
Table 12: The Prompt Template for Knowledge State Query on MMLU. The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
In-Context Examples Answer the following question as simple as possible. Question: {Question1} Answer: {Answer1} Answer the following question as simple as possible. Question: {Question2} Answer: {Answer2} Answer the following question as simple as possible. Question: {Question3} Answer: {Answer3}
Instruction Answer the following question as simple as possible. Question: {Question} Answer:
Table 13: The Prompt Template for Knowledge State Query on TriviaQA. The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
Prompts for training.
For the rehearsal training and Van-Tuning, we use the basic prompt as shown in Table 12 and Table 13 without in-context example. All other experiments use the refuse prompt as shown in Table 14 and Table 15. Loss is only computed on the target answer $\{\text{answer}_{rait}\}$.
Instruction There is a single choice question about {Task}. If you know the answer, please directly respond with the correct answer A, B, C, or D. If you do not know the answer, please respond with “I don’t know.”. Question:{Question} A. {Content_of_A} B. {Content_of_B} C. {Content_of_C} D. {Content_of_D} Answer: {$Answer_{\text{rait}}$}
Table 14: The REFUSE Prompt Template for Training on MMLU. The The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
Instruction Answer the following question, and if you don’t know the answer, only reply with “I don’t know”:{Question} {$Answer_{\text{rait}}$}
Table 15: The REFUSE Prompt Template for Training on TriviaQA. The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
Prompts for evaluation.
The Init-Basic method uses the original question format for evaluation, without any prior instructions. For the other methods, the evaluation prompts are shown in Tables 16 and 17.
Instruction There is a single choice question about {Task}. If you know the answer, please directly respond with the correct answer A, B, C, or D. If you do not know the answer, please respond with “I don’t know.”. Question:{Question} A. {Content_of_A} B. {Content_of_B} C. {Content_of_C} D. {Content_of_D} Answer:
Table 16: The REFUSE Prompt Template for Evaluation on MMLU. The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
Instruction Answer the following question, and if you don’t know the answer, only reply with “I don’t know”: {Question}
Table 17: The REFUSE Prompt Template for Evaluation on TriviaQA. The Italic {text} in Curly Braces Represents Variables That Need To be Replaced.
D.3 Training
LoRA or full parameter tuning
In the MCQA task, we used LoRA, setting $r=64$, $\alpha=16$, $dropout=0.1$. In the OEQA task, we used full parameter training.
Learning rate
We use the cosine learning rate scheduler with warming up. The maximum learning rate for MCQA task is 2e-4. The maximum learning rate on OEQA for LLaMA2-7B-Chat is 2e-5, for LLaMA3-8B-Instruct is 8e-6.
D.4 Evaluation
During the evaluation, the LLM answers questions in the validation set in generation mode. When parsing the results, the algorithm first checks whether the LLM’s response is refusal reply. We have constructed a list of refusal keywords, and the algorithm checks whether the response generated by the LLM contains any refusal keywords. If so, it is considered that the model refuses to answer the question; otherwise, it is considered that the LLM has answered normally. Conversely, for questions that the LLM answers normally, we parse the answers by rules and calculate metrics based on the ground truth. Specifically, for the MCQA tasks (MMLU, ARC-c), we parse the LLM’s response by rules to obtain the four options ABCD; for the OEQA tasks (TriviaQA, NQ), the answers in the dataset usually contain one or more candidates. If any candidate is included in the LLM’s response, it is considered that the LLM has answered correctly; otherwise, it is incorrect.

### v2

@@
 However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal.
-In this paper, we explore two primary causes of over-refusal: Static conflict emerges when the RAIT data is constructed solely on correctness criteria, causing similar samples in the LLM’s feature space to be assigned different labels (original vs.
+In this paper, we explore two primary causes of over-refusal: Static conflict occurs when similar samples within the LLM’s feature space receive differing supervision signals (original vs.
 modified “I don’t know”).
-Dynamic conflict occurs due to the changes of LLM’s knowledge state during fine-tuning, which transforms previous unknown questions into knowns, while the training data, which is constructed based on the initial LLM, remains unchanged.
+Dynamic conflict, on the other hand, emerges as the LLM’s knowledge evolves during SFT, allowing it to answer questions that were previously unanswerable.
+Yet, these now-answerable training samples still retain the original “I don’t know” supervision signals based on the initial LLM state, resulting in inconsistencies.
 These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal.
-To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT).
+To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Tuning (CRaFT).
 CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts.
@@
 Finally, we fine-tune the initial LLM using the acquired RATI data, achieving robust overall performance.
+In conducting our experimental analysis, we sought a well-founded metric within current research.
+However, existing methods have notable limitations.
+For instance, (Feng et al.
+2024) proposed using multiple evaluation metrics simultaneously, which is challenging to manage as it is difficult for a model to optimize across multiple criteria at once.
+Additionally, we provided counterexamples to demonstrate the inadequacy of metrics proposed in other studies.
+Consequently, we examined these shortcomings and introduced a singular and comprehensive metric: Truthful Helpfulness Score(THS).
 Overall, our main contributions are as follows:
@@
 We term this “static conflict” to distinguish it from another conflict type discussed later.
-Figure 4: RAIT Data Construction.
+Figure 4: RAIT Data Construction of Cor-RAIT, CorCer-RAIT and CRaFT.
+Cor-RAIT partitions data based on accuracy $\mu$ and a threshold $\tau_{\mu}$.
+For CorCer-RAIT, $D_{\text{van}}$ is derived from samples with accuracy exceeding the threshold and the highest certainty, while $D_{\text{idk}}$ consists of samples with accuracy below the threshold and the lowest certainty.
+CRaFT employs a two-stage process: in the first stage, data with $\Delta\mu>0$ is excluded through knowledge queries; the second stage follows the same procedure as CorCer-RAIT.
 3.3 Certainty Mitigates the Static Conflicts
-Theoretical analysis (see Appendix A.2) indicates a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$.
+We conducted a theoretical analysis 333detailed proof in Appendix A.2 establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$.
 This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively.
@@
 We suggest adopting the certainty of the LLM’s response as the indicator variable.
-Our theoretical analysis (see Appendix A.3) shows that certainty meets the above requirements.
+Our theoretical analysis shows that certainty meets the above requirements.
 [^4] Section 4.2 discusses various methods for representing LLM response certainty.
@@
 We term it as “dynamic conflict” of RAIT data, which significantly contributes to the over-refusal in Cor-RAIT.
-We use rehearsal training (see Chapter 4.3) to capture the LLM’s natural knowledge flow during SFT.
-The experiment on MMLU dataset (Hendrycks et al.
-2020) and LLaMA-3-8B-instruct showed that the correctness of 69% samples, initially below 0.5, improved, thereby validating the above analysis on dynamic conflict.
-More experimental results are in Appendix A.5.
+We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2.
+Rehearsal training is designed to capture the LLM’s natural knowledge flow during SFT.
+Experiments on the MMLU dataset (Hendrycks et al.
+2020) and LLaMA-3-8B-instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict.
+Additional experimental results are provided in Appendix A.5.
 4 Methodology
@@
 CRaFT contains two stages, as shown in Figure 6.
-Figure 6: The Framework of CRaFT.
+Figure 6: The Framework of CRaFT: Stage 1 queries the knowledge state and flow, while Stage 2 constructs and tunes refusal-aware instructions.
 Stage 1: Query the Knowledge State and Flow of LLM
-[…]
+The output of stage one is the knowledge state and flow indicators of the model.
+First, we perform a knowledge state query to obtain the correctness and certainty of the model’s responses to the samples in the source dataset.
+Next, we conduct rehearsal training on the model, resulting in the perturbed version.
+By comparing the knowledge states before and after perturbation, we derive the indicators of knowledge flow during the supervised fine-tuning process.
 Stage 2: Refusal-Aware Instructions Construction and Tuning
-Using the knowledge state and flow from Stage 1, we select suitable samples from $D_{\text{src}}$ to construct the RAIT data $D_{\text{RAIT}}$ (see Section 4.4), which is is used to fine-tune the initial model $\mathcal{M}_{0}$.
-4.2 Knowledge State Query
+Using the knowledge state and flow from Stage 1, we select suitable samples from $D_{\text{src}}$ to construct the RAIT data, which is is used to fine-tune the initial model.
+4.2 Query the Knowledge State and Flow of LLM
+Knowledge State Query
 […]
@@
 More details about knowledge state query are in Appendix B.1.
-4.3 Rehearsal Training and Knowledge Flow
+Rehearsal Training and Knowledge Flow
 […]

### v3

@@
 modified “I don’t know”).
-Dynamic conflict, on the other hand, emerges as the LLM’s knowledge evolves during SFT, allowing it to answer questions that were previously unanswerable.
-Yet, these now-answerable training samples still retain the original “I don’t know” supervision signals based on the initial LLM state, resulting in inconsistencies.
+Dynamic conflict arises as the LLM’s evolving knowledge during SFT enables it to answer previously unanswerable questions, but the now-answerable training samples still retain the original “I don’t know” supervision signals from the initial LLM state, leading to inconsistencies.
 These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal.
@@
 Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process.
-Source code and training data will be released at Github.
+Code & Datahttps://github.com/opendatalab/CRaFT
 1 Introduction
@@
 2023; Dubey et al.
-2024; Yang et al.
+2024; Touvron et al.
+2023; Yang et al.
 2024).
@@
 Ideally, the reliable LLM should actively refuse to answer questions it doesn’t know to avoid incorrect responses (Wen et al.
-; Li et al.
+2024; Li et al.
 2024b).
@@
 2024; Xu et al.
-2024; Cheng et al.
+2024a; Cheng et al.
+2024; Xu et al.
+2024a; Bai et al.
+2023; Cheng et al.
 2024) that Refusal-Aware Instruction Tuning (RAIT) can enable LLMs to refuse answering questions beyond their knowledge.
@@
 Generally, the RAIT process cab be described as follows: the initial LLM is prompted to answer all questions in the train set $D_{\text{src}}$.
-Based on the response correctness, the samples are categorized into two groups.
-Samples with correct responses are considered as known knowledge of the LLM.
-Their answers will not be modified, referred to as vanilla samples $D_{\text{van}}$.
-Conversely, samples with incorrect responses are considered as unknown knowledge of LLM.
-Their original answers are replaced with refusal responses like “I don’t know”.
-These samples are called IdK samples $D_{\text{idk}}$.
-The constructed RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$ is used to fine-tune the initial LLM, improving its ability to refuse to answer unknown questions.
+Based on the correctness of the responses, the samples are divided into two groups.
+Correct responses are considered known knowledge and are labeled as vanilla samples $D_{\text{van}}$, whose answers remain unchanged.
+Incorrect responses are considered unknown knowledge and are replaced with ”I don’t know”, forming the IdK samples $D_{\text{idk}}$.
+The combined RAIT data, $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$, is used to fine-tune the LLM, enhancing its ability to refuse unknown questions.
 For simplicity, we refer to this original form of the RAIT method as Cor-RAIT.
-However, (Cheng et al.
-2024) reveals that the Cor-RAIT causes the fine-tuned LLM to refuse some questions that could have been answered correctly.
-Experiments demonstrate significant drop of LLM’s accuracy after Cor-RAIT: on the TriviaQA dataset (Joshi et al.
-2017), it decreases from 45.05% to 28.57%.
-Similarly, on the Natural Questions dataset (Kwiatkowski et al.
-2019), accuracy fell from 24.65% to 15.93%.
-We refer to this phenomenon as over-refusal, as shown in Figure 1.
 Figure 2: Two causes of over-refusal: (a) Static conflict means the similar samples in the LLM’s feature space being assigned different labels (original vs.
@@
 These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal.
+However, (Cheng et al.
+2024) shows that Cor-RAIT causes the fine-tuned LLM to refuse some questions that could have been answered correctly.
+Experiments reveal a significant accuracy drop after Cor-RAIT: on the TriviaQA dataset (Joshi et al.
+2017), accuracy falls from 45.05% to 28.57%, and on the Natural Questions dataset (Kwiatkowski et al.
+2019), it drops from 24.65% to 15.93%.
+We refer to this phenomenon as over-refusal, as shown in Figure 1.
 In addressing the over-refusal brought by Cor-RAIT, we identified two primary causes as shown in Figure 2.
-(1) Static Conflict: In the LLM feature space, two closely located samples might be assigned to $D_{\text{van}}$ and $D_{\text{idk}}$ under the Cor-RAIT framework.
+(1) Static Conflict: In the LLM representation space, two closely located samples might be assigned to $D_{\text{van}}$ and $D_{\text{idk}}$ under the Cor-RAIT framework.
 As illustrated by t-SNE in Figure 3(a), significant intersections exist between $D_{\text{van}}$ and $D_{\text{idk}}$, complicating their differentiation.
@@
 Research (Ren et al.
-2024; Ren and Sutherland 2024) shows that the knowledge state of LLMs changes during Supervised Fine-Tuning (SFT), with questions potentially shifting from unknown to known and vice versa.
+2024; Ren and Sutherland 2024; Xu et al.
+2024b) shows that the knowledge state of LLMs changes during Supervised Fine-Tuning (SFT), with questions potentially shifting from unknown to known and vice versa.
 This phenomenon is reminiscent of Heraclitus’ saying, “no man ever steps in the same river twice.” However, current methods use static RAIT data reflecting the initial LLM’s knowledge state throughout SFT, which ignores these changes.
-This oversight leads to conflicts between the RAIT data and the LLM’s evolving knowledge, resulting in nefficient training and over-refusal.
+This oversight leads to conflicts between the RAIT data and the LLM’s evolving knowledge, resulting inefficient training and over-refusal.
 To address the two problems above, we propose Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT).
 Our approach consists of two stages.
-Stage 1: Query the knowledge state and flow of LLM.
-First, the knowledge state of the initial LLM is probed.
-Compared to Cor-RAIT, we introduce response certainty on top of the LLM’s response correctness.
-The joint use of these two indicator variables effectively alleviates the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$.
-Further, to capture the dynamic changes in LLM’s knowledge state during training, we introduce a rehearsal training mechanism.
-This mechanism fine-tunes the initial LLM using data samples highly consistent with LLM’s internal knowledge.
-By not introducing any new knowledge during training(Ren et al.
+Stage 1: Querying the Knowledge State and Flow of the LLM.
+First, we probe the initial LLM’s knowledge state.
+Unlike Cor-RAIT, we incorporate response certainty alongside correctness, effectively alleviating the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$.
+To capture the LLM’s dynamic knowledge changes during training, we introduce a rehearsal training mechanism.
+This fine-tunes the LLM with data samples that align closely with its internal knowledge, without introducing new knowledge (Ren et al.
 2024; Kang et al.
-2024), this method enables us to observe the LLM’s natural internal adjustments to its knowledge state.
-The difference in knowledge states between the fine-tuned and the initial LLMs helps characterize the knowledge flow within the LLM during training, aiding in the identification and resolution of dynamic conflicts.
+2024).
+This approach allows us to observe the LLM’s natural knowledge adjustments.
+The differences between the fine-tuned and initial LLMs reveal the knowledge flow during training, helping to identify and resolve dynamic conflicts.
 Stage 2: Refusal-aware instructions construction and tuning.
-Taking into account both the static knowledge state and dynamic knowledge flow, we filter out vanilla samples and IdK samples from the original RAIT data, therefore effectively reduce both of the static and dynamic conflicts.
-Finally, we fine-tune the initial LLM using the acquired RATI data, achieving robust overall performance.
+By considering both the static knowledge state and dynamic knowledge flow, we filter out vanilla and IdK samples from RAIT data, reducing conflicts.
+We then fine-tune the initial LLM with the refined data, improving overall performance.
 In conducting our experimental analysis, we sought a well-founded metric within current research.
-However, existing methods have notable limitations.
-For instance, (Feng et al.
-2024) proposed using multiple evaluation metrics simultaneously, which is challenging to manage as it is difficult for a model to optimize across multiple criteria at once.
-Additionally, we provided counterexamples to demonstrate the inadequacy of metrics proposed in other studies.
+Existing methods have notable limitations, either proposing multiple metrics that are hard to optimize simultaneously or relying on inherently flawed metrics, as demonstrated by our counterexamples.
 Consequently, we examined these shortcomings and introduced a singular and comprehensive metric: Truthful Helpfulness Score(THS).
@@
   • We conducted the in-depth analysis of static and dynamic conflicts in existing correctness-based RAIT data, revealing that they cause the trained LLMs’ mis-classification of known and unknown questions, leading to the issue of over-refusal in current RAIT methods.
-  • To address these static and dynamic conflicts, we introduce CRaFT: during the construction of RAIT data, instead of solely depending on correctness, we additionally incorporate certainty to reduce static conflicts; through rehearsal training, we capture the trends of the LLM’s knowledge flow, which helps effectively mitigate dynamic conflicts.
-Extensive experiments show that CRaFT can effectively alleviate over-refusal and enhance the overall performance of the LLMs.
-  • Given the shortcomings of existing refusal-aware metrics, we developed the Truthful Helpfulness Score (THS).
-This metric balances the reliability and helpfulness of LLMs, enabling a thorough and precise evaluation of their overall performance.
+  • To address static and dynamic conflicts, we introduce CRaFT: it reduces static conflicts by incorporating certainty alongside correctness during RAIT data construction, and mitigates dynamic conflicts through rehearsal training to capture knowledge flow trends.
+Extensive experiments demonstrate that CRaFT alleviates over-refusal and improves overall LLM performance.
+  • We analyze the shortcomings of existing refusal-aware metrics and introduce the Truthful Helpfulness Score (THS), which balances reliability and helpfulness for a comprehensive evaluation of LLM performance.
     Figure 3: t-SNE visualization of the LLM feature space
@@
 2.1 Mitigating Hallucinations of LLMs
-Researchers have developed several methods to mitigate hallucinations of LLMs, including data augmentation(Neeman et al.
-2022), improving decoding strategies(Holtzman et al.
+Researchers have developed various methods to mitigate LLM hallucinations, including data augmentation (Neeman et al.
+2022), improved decoding strategies (Holtzman et al.
 2019; Chuang et al.
-2023), integrating external knowledge (Karpukhin et al.
+2023), external knowledge integration (Karpukhin et al.
 2020), knowledge editing (Zhang, Yu, and Feng 2024; Li et al.
-2024a) and honesty alignment(Zhang et al.
+2024a), and honesty alignment (Zhang et al.
 2024; Xu et al.
+2024a; Bai et al.
 2024).
-Unlike the previous approach of correcting model answers, honesty alignment encourages models to say “I don’t know” when answering unknown questions.
-Ways to achieve honesty alignment include refusal-aware instruction tuning (RAIT) and reinforcement learning-based honesty alignment (RLHA).
+Unlike traditional correction methods, honesty alignment encourages models to say “I don’t know” for unknown questions.
 2.2 Refusal-Aware Instruction Tuning
@@
 R-Tuning (Zhang et al.
-2024) identifies these unknown questions by having the LLM answer each question in the training set a single time and then verifying whether the LLM’s responses match the label.
+2024) identifies these questions by having the LLM answer each once and verifying response accuracy.
 In (Yang et al.
-2023), the LLM is prompted to answer the same question multiple times.
-The target answer is then modified based on the correctness ratio of the responses.
-In (Wan et al.
-2024), a knowledge-based verification mechanism ensures the model’s knowledge to be consistent with external trusted sources, increasing refusal accuracy and preventing the spread of false information.
-2.3 Reinforcement Learning-based Honesty Alignment
-RLHA improves model refusal ability by reinforcement learning with well-designed reward models.
-(Xu et al.
-2024) proposed a novel alignment framework called RLKF, which used knowledge feedback to dynamically determine the knowledge boundary of LLM and train reliable reward models to encourage refusal responses.
-(Cheng et al.
-2024) introduces a refusal function that learns whether the model should refuse to answer by evaluating the extent to which the input aligns with human honesty.
+2023), the LLM answers the same question multiple times, with the target answer adjusted based on the correctness ratio.
+(Wan et al.
+2024) uses a knowledge-based verification mechanism to ensure consistency with trusted external sources, enhancing refusal accuracy and preventing misinformation.
 3 Over-Refusal: Analysis and Insights
 3.1 Refusal-Aware Instruction Tuning
-Given the initial LLM $\mathcal{M}_{0}$ and the instruction dataset $D_{\text{src}}$ consisting of question and answer pairs $x=(q,a)$, we select and modify data from $D_{\text{src}}$ to construct the data $D_{\text{rait}}$, composed of pairs $(q,a_{\text{rait}})$.
-$D_{\text{rait}}$ is then used to perform SFT on $\mathcal{M}_{0}$, resulting in the new LLM with the capability to decline unknown questions.
-This process is referred to as Refusal-Aware Instruction Tuning (RAIT).
+Given the initial LLM $\mathcal{M}_{0}$ and the instruction dataset $D_{\text{src}}$ of question-answer pairs $x=(q,a)$, we modify $D_{\text{src}}$ to construct $D_{\text{rait}}$, consisting of pairs $(q,a_{\text{rait}})$.
+$D_{\text{rait}}$ is then used for SFT on $\mathcal{M}_{0}$, resulting in a new LLM capable of declining unknown questions, a process called Refusal-Aware Instruction Tuning (RAIT).
 Existing studies (Zhang et al.
@@
 2023; Cheng et al.
-2024) use $\mathcal{M}_{0}$ infer questions from $D_{\text{src}}$ and assess their correctness, denoted as $\mu$.
-As illustrated in Figure 4(a), the correctness threshold $\tau_{\mu}$ is defined firstly.
-Then, if $\mu<\tau_{\mu}$, the answer for the question is transformed into a new target answer, “I don’t know” and assigned to the subset $D_{\text{idk}}$, comprising what are termed as IdK samples.
-Conversely, if $\mu\geq\tau_{\mu}$, the original answer remains unchanged and is assigned to the subset $D_{\text{van}}$, which includes the so-called vanilla samples.
-The resulting RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$.
-This Correctness-based RAIT is referred to as Cor-RAIT.
+2024) use $\mathcal{M}_{0}$ to infer and assess the correctness of questions in $D_{\text{src}}$, denoted as $\mu$.
+As shown in Figure 4(a), the correctness threshold $\tau_{\mu}$ is first defined.
+If $\mu<\tau_{\mu}$, the answer is changed to ”I don’t know” and assigned to the IdK subset $D_{\text{idk}}$.
+If $\mu\geq\tau_{\mu}$, the original answer remains, and the pair is assigned to the vanilla subset $D_{\text{van}}$.
+The resulting RAIT data is $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$, and this correctness-based RAIT is called Cor-RAIT.
 However, LLMs exhibited significant over-refusal after Cor-RAIT, as shown in Figure 1. Subsequent sections of this chapter will analyze the causes and offer practical insights.
 3.2 Static Conflicts in Cor-RAIT
-During the RAIT process, LLMs learn to reject unknown samples by supervisions from $D_{\text{idk}}$.
-Our insight is that if the Cor-RAIT dataset contains vanilla and IdK samples that are closely positioned in the LLM’s representation space, the trained LLM may mistakenly classify similar vanilla samples (known questions) as IdK samples (unknown questions), causing over-refusals.
+During the Cor-RAIT process, LLMs learn to reject unknown samples by supervisions from $D_{\text{idk}}$.
+Our insight is that if the Cor-RAIT dataset contains vanilla and IdK samples that are closely positioned in the LLM’s representation space, the trained LLM may mistakenly classify similar vanilla samples as IdK samples, causing over-refusals.
 To verify this, we analyzed the sample distributions of $D_{\text{van}}$ and $D_{\text{idk}}$.
@@
 2024) representation space, where IdK and vanilla samples have significant intersections.
-[^1] We combined additional datasets and LLMs; see Appendix A.1 for similar observations.
+[^1] More experiments involving additional datasets and LLMs are provided in Appendix A.1.
 Furthermore, we introduce the Conflict Rate for Similar Samples (CRSS) to quantitatively assess conflicts in supervision signals among similar samples in the RAIT dataset.
@@
 The above analysis reveals that Cor-RAIT generates numerous similar sample pairs between $D_{\text{van}}$ and $D_{\text{idk}}$, resulting in conflicting supervision signals which leads to over-refusal.
-We term this “static conflict” to distinguish it from another conflict type discussed later.
+We term this static conflict to distinguish it from another conflict type discussed later.
 Figure 4: RAIT Data Construction of Cor-RAIT, CorCer-RAIT and CRaFT.
@@
 3.3 Certainty Mitigates the Static Conflicts
-We conducted a theoretical analysis 333detailed proof in Appendix A.2 establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$.
+We conducted a theoretical analysis 333detailed proof in Appendix A.2.
+establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$.
 This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively.
@@
 This variable should ensure that similar samples share comparable values, reducing the above misclassification.
-We suggest adopting the certainty of the LLM’s response as the indicator variable.
+We suggest adopting the certainty (Jiang et al.
+2023) of the LLM’s response as the indicator variable.
 Our theoretical analysis shows that certainty meets the above requirements.
@@
 The phenomenon, which we refer to as “knowledge flow”, can cause previously incorrectly answered questions to become correct ones during SFT.
-Despite this dynamic evolutions, the target answer of the training data remains static during the SFT process, which reflects the knowledge state of the initial LLM but ignores subsequent changes.
-We term it as “dynamic conflict” of RAIT data, which significantly contributes to the over-refusal in Cor-RAIT.
-We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2.
+Despite this dynamic evolutions, the target answer of the training data remains static during the RAIt, which reflects the knowledge state of the initial LLM but ignores subsequent changes.
+We term it as dynamic conflict, which significantly contributes to the over-refusal in Cor-RAIT.
+We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2..
 Rehearsal training is designed to capture the LLM’s natural knowledge flow during SFT.
 Experiments on the MMLU dataset (Hendrycks et al.
-2020) and LLaMA-3-8B-instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict.
+2020) and LLaMA3-8B-Instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict.
 Additional experimental results are provided in Appendix A.5.
@@
 CRaFT contains two stages, as shown in Figure 6.
-Figure 6: The Framework of CRaFT: Stage 1 queries the knowledge state and flow, while Stage 2 constructs and tunes refusal-aware instructions.
+Figure 6: The Framework of CRaFT: Stage 1 queries knowledge state and flow, while Stage 2 constructs RAI data and tunes.
 Stage 1: Query the Knowledge State and Flow of LLM
@@
 Correctness is obtained through exact match across the $N$ responses, calculating the proportion of accurate answers.
-Certainty is evaluated using a pretrained SentenceTransformer model  [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode each response $\hat{a}_{j}$ into its vector representation $E(\hat{a}_{j})$, and the average similarity is computed between these vector embeddings (excluding diagonal elements).
+Certainty is evaluated using a pretrained SentenceTransformer model  [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode each response $\hat{a}_{j}$ into embedding $E(\hat{a}_{j})$, and the average similarity is computed between these embeddings (excluding diagonal elements).
 The correctness values range from $[0,1]$.
-In MCQA task, certainty ranges from [$-Entropy(N_{o})$, 0], and for OEQA, from $[0,1]$.
-$N_{o}$ stands for the number of options.
+In MCQA task, certainty ranges from [$-log|O|$, 0], and for OEQA, from $[0,1]$.
 More details about knowledge state query are in Appendix B.1.
@@
 More details about rehearsal training are in Appendix B.2.
-4.3 Refusal-Aware Instruct Data Constuction
-In contrast to Cor-RAIT, which selects RAIT samples solely based on correctness, our approach leverages four parameters $\mu$, $\sigma$, $\Delta\mu$, and $\Delta\sigma$ to characterize both the knowledge state and flow of $\mathcal{M}_{0}$.
+4.3 Refusal-Aware Instructions Constuction and Tuning
+Unlike Cor-RAIT, which selects RAIT samples solely based on correctness, our approach leverages four parameters $\mu$, $\sigma$, $\Delta\mu$, and $\Delta\sigma$ to characterize both the knowledge state and flow of $\mathcal{M}_{0}$.
 The challenge lies in making informed sample selections across these four dimensions.
@@
 5.1 Dataset
-In this paper, we evaluate two tasks: knowledge-oriented Multiple Choice Questions Answer (MCQA) and Open-ended Questions Answer (OEQA).
-For the MCQA task, the MMLU (Hendrycks et al.
-2020) test split is used as the training set, the MMLU val split is used as the In-Domain (ID) test set, and the ARC-c (Clark et al.
-2018) test split is the Out-Of-Domain (OOD) test set.
-For the OEQA task, the TriviaQA (Joshi et al.
-2017) train split is used as the training set, the TriviaQA dev split is the ID test set, and the NQ (Kwiatkowski et al.
-2019) dev split is the OOD test set.
-Refer to Appendix D.1 for details of the dataset.
+We evaluate two tasks: knowledge-oriented Multiple Choice Questions Answering (MCQA) and Open-ended Questions Answering (OEQA).
+For MCQA, the MMLU (Hendrycks et al.
+2020) test split serves as the training set, MMLU val as the In-Domain (ID) test set, and ARC-c (Clark et al.
+2018) test split as the Out-Of-Domain (OOD) test set.
+For OEQA, the TriviaQA (Joshi et al.
+2017) train split is used for training, TriviaQA dev for the ID test set, and NQ (Kwiatkowski et al.
+2019) dev for the OOD test set.
+More details are in Appendix D.1.
 5.2 Metric
-In the post-RAIT evaluation of LLMs, each test sample is classified as correct, incorrect, or refused.
-We calculate the accuracy ($P_{c}$), error ($P_{w}$), and refusal rates ($P_{r}$) based on these categories.
-These metrics reveal trade-offs in LLM performance, prompting the essential question: How to determine the best-performing model?
+In post RAIT evaluation of LLMs, each test sample is classified as correct, incorrect, or refused.
+We calculate accuracy ($P_{c}$), error ($P_{w}$), and refusal rates ($P_{r}$) to assess performance, highlighting the key question: How to identify the better-performing model?
 Shortcomings of existing refusal-aware metrics
@@
 TRUTHFUL (Cheng et al. 2024) $\uparrow$ (0.8) (0.75) 1 1
-rely (Xu et al. 2024) $\uparrow$ (0.55) (0.548) 0.75 1
+rely (Xu et al. 2024a) $\uparrow$ (0.55) (0.548) 0.75 1
 R-Acc (Feng et al. 2024) $\uparrow$ (0.8) (0.778) 1 1
@@
 However, existing metrics exhibit significant issues, as indicated by the numbers in (parentheses).
-We constructed the initial model $\mathcal{M}_{0}$ along with four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$, with exhibiting exemplary properties.
-Performance comparisons clearly demonstrate a progressive improvement: $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$.
-Detailed information on $\mathcal{M}_{0}$ through $\mathcal{M}_{4}$ can be found in Appendix C.2.
-However, existing metrics exhibit significant issues: $S_{\text{honesty}}$ (Yang et al.
-2023) inaccurately rates $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$ and equates $\mathcal{M}_{3}$ with $\mathcal{M}_{4}$.
-TRUTHFUL (Cheng et al.
-2024) mistakenly favor $\mathcal{M}_{1}$ over $\mathcal{M}_{2}$.
-R-Acc, ER, A-Acc and A-F1 (Feng et al.
-2024) also incorrectly ranks $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$.
-AP (Zhang et al.
-2024) fails to differentiate between $\mathcal{M}_{3}$ and $\mathcal{M}_{4}$.
+We constructed an initial model $\mathcal{M}_{0}$ and four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$, showing progressive improvement: $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$.
+Details on these models are in Appendix C.2.
+However, existing metrics have notable flaws: $S_{\text{honesty}}$ (Yang et al.
+2023) ranks $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$ and treats $\mathcal{M}_{3}$ the same as $\mathcal{M}_{4}$; TRUTHFUL (Cheng et al.
+2024) favors $\mathcal{M}_{1}$ over $\mathcal{M}_{2}$; and R-Acc, ER, A-Acc, and A-F1 (Feng et al.
+2024) also rank $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$.
+Additionally, AP (Zhang et al.
+2024) fails to distinguish between $\mathcal{M}_{3}$ and $\mathcal{M}_{4}$.
 Figure 7: Truthful Helpfulness Score (THS).
@@
 5.3 Baselines
-To verify the effectiveness of CRaFT, we compared it with the current mainstream methods: Init-Basic: We use the initial LLM, based on common question-answering prompts to instruct the LLM to answer questions.
-Init-Refuse: Compared to Init-Basic, we add instructions like “If you do not know the answer, please respond with ‘I don’t know.’” in the prompt.
-Van-Tuning: We randomly select $N_{\text{van}}+N_{\text{idk}}$ samples from $D_{\text{src}}$ for instruct-tuning directly, without sample modification.
-Cor-RAIT: We implemented this method following the settings in (Zhang et al.
-2024), which is based on the correctness of the reply to filter and modify RAIT data.
-The prompts involved in each baseline are detailed in the Appendix D.2.
+To verify CRaFT’s effectiveness, we compared it with mainstream methods: Init-Basic: Uses the initial LLM with common question-answering prompts.
+Init-Refuse: Adds instructions like “If you don’t know, respond with ‘I don’t know.’”.
+Van-Tuning: Randomly selects $N_{\text{van}}+N_{\text{idk}}$ samples from $D_{\text{src}}$ for instruct-tuning without modification.
+Cor-RAIT: Implements the method from (Zhang et al.
+2024), filtering and modifying RAIT data based on response correctness.
+Detailed prompts for each baseline are in Appendix D.2.
 5.4 Implementation Details
-In the experiments, we selected LLaMA2-7B-Chat (Touvron et al.
+In the experiments, we used LLaMA2-7B-Chat (Touvron et al.
 2023) and LLaMA3-8B-Instruct (Dubey et al.
-2024) as the the initial LLM $\mathcal{M}_{0}$.
-In the MCQA task, we selected 5000 samples from MMLU as training data; in the OEQA tasks, we selected 10,000 samples from TriviaQA as training data.
-Except for Van-Tuning, where all samples are vanilla samples, in other RAIT settings, the ratio of vanilla and IdK samples is 1:4.
-In MCQA and OEQA tasks 5-shot and 3-shot were used in knowledge state query respectively.
-More details of the LLM’s knowledge state and flow query are in Appendix B.
-We used XTuner  [^7] https://github.com/InternLM/xtuner for all Instruct Tuning experiments, with 3 epochs and a maximum context length of 2048.
-In the MCQA tasks, we used LoRA, setting $r=64$, $\alpha=16$, $dropout=0.1$, and a learning rate of 2e-4; in the OEQA tasks, we used full parameter training.
-More details of training are in Appendix D.3.
-0-shot and greedy decoding setting were used in evaluation.
-More details of evaluation are in Appendix D.4.
-OpenCompass  [^8] https://github.com/open-compass/opencompass was used for all knowledge state queries and evaluations.
-All experiments were conducted on NVIDIA A100-80GB GPUs.
+2024) as the initial LLM $\mathcal{M}_{0}$.
+For the MCQA task, we selected 5000 samples from MMLU, and for the OEQA task, we used 10,000 samples from TriviaQA as training data.
+In all RAIT settings, except Van-Tuning, the ratio of vanilla to IdK samples was 1:4.
+We applied 5-shot and 3-shot knowledge state queries for the MCQA and OEQA tasks, respectively.
+Details on knowledge state and flow queries are in Appendix B.
+For Instruct Tuning, we used XTuner  [^7] https://github.com/InternLM/xtuner with 3 epochs and a maximum context length of 2048.
+In MCQA, we applied LoRA (Hu et al.
+2021) with settings $r=64$, $\alpha=16$, dropout=0.1, and a learning rate of 2e-4; for OEQA, full parameter training was used.
+More details on training are in Appendix D.3.
+We used 0-shot and greedy decoding for evaluation, with further details in Appendix D.4.
+OpenCompass  [^8] https://github.com/open-compass/opencompass was employed for knowledge state queries and evaluations.
+All experiments were run on NVIDIA A100-80GB GPUs.
 6 Experimental Results and Analyses
+LLMs QA Type MCQA OEQA
+Dataset MMLU (ID) ARC-c (OOD) TriviaQA (ID) NQ (OOD)
+[…]
+LLaMA2-7B Chat Baselines Init-Basic 45.6 52.8 00.0 53.9 46.0 00.0 54.0 46.0 00.0 29.3 70.7 00.0
+Init-Refuse 36.4 38.9 03.9 44.4 35.7 02.6 37.0 21.7 11.5 20.8 38.6 04.8
+Van-Tuning 46.9 53.1 01.2 54.5 45.5 01.2 48.6 44.5 -03.7 18.3 50.2 -02.5
+Cor-RAIT 44.5 39.6 11.3 55.8 38.1 11.1 41.3 18.3 19.7 16.2 27.6 04.7
+[…]
+[…]
+w/o Cer 38.4 32.1 11.5 52.5 32.9 13.9 38.5 15.7 20.1 14.6 22.1 05.4
+LLaMA3-8B Instruct Baselines Init-Basic 66.8 33.1 00.0 80.6 19.5 00.0 66.8 33.2 00.0 40.3 59.7 00.0
+Init-Refuse 50.0 17.0 15.6 65.3 14.4 05.6 53.9 20.8 12.0 31.1 38.6 05.0
+Van-Tuning 69.5 30.5 08.0 80.3 19.7 -01.3 55.0 38.1 -21.8 21.0 48.5 -11.7
+Cor-RAIT 63.9 21.6 20.4 79.4 16.2 12.2 45.4 13.2 18.8 17.2 25.6 -00.1
+[…]
+[…]
+w/o Cer 62.1 18.4 25.0 78.2 17.3 06.5 43.0 11.2 20.5 15.8 23.5 -00.1
+Table 2: Performance comparisons on MMLU, ARC-c, TriviaQA and NQ. The best performance is highlighted in boldface.
 6.1 Overall Performance
-The experimental results on the OEQA and MCQA tasks are presented in Table 2 and 3.
+The experimental results on the OEQA and MCQA tasks are presented in Table 2.
 Under the ID setting for both types of tasks, our method outperformed four baseline models on THS, achieving the best results.
@@
 Therefore, although CRaFT surpassed Cor-RAIT under all tasks and settings, the improvement was limited under the OOD setting for OEQA due to training paradigm.
-TriviaQA (ID) NQ (OOD)
-[…]
-Llama2- 7B- Chat Init-Basic 54 46 0 29.3 70.7 0
-Init-Refuse 37 21.7 11.5 20.8 38.6 4.8
-Van-Tuning 48.6 44.5 -3.7 18.3 50.2 -2.5
-Cor-RAIT 41.3 18.3 19.7 16.2 27.6 4.7
-CRaFT 38.5 12.9 23.3 15.8 22.4 6.5
-Llama3- 8B- Instruct Init-Basic 66.8 33.2 0 40.3 59.7 0
-Init-Refuse 53.9 20.8 12 31.1 38.6 5
-Van-Tuning 55 38.1 -21.9 21 48.5 -11.7
-Cor-RAIT 45.4 13.2 18.8 17.2 25.6 -0.1
-CRaFT 43.5 10.9 21.5 19 27.5 0.4
-Table 2: Experimental Results in OEQA Task.
-MMLU (ID) ARC-c (OOD)
-[…]
-Llama2- 7B- Chat Init-Basic 45.6 52.8 0 53.9 46 0
-Init-Refuse 36.4 38.9 3.9 44.4 35.7 2.6
-Van-Tuning 47 53.1 1.2 54.5 45.5 1.2
-Cor-RAIT 44.5 39.6 11.3 55.8 38.1 11.1
-CRaFT 43.9 36.4 12.5 54.7 35.9 12.6
-Llama3- 8B- Instruct Init-Basic 66.8 33.1 0 80.6 19.5 0
-Init-Refuse 50 17 15.7 65.3 14.4 5.6
-Van-Tuning 69.5 30.5 8 80.3 19.7 -1.3
-Cor-RAIT 63.9 21.6 20.4 79.4 16.2 12.2
-CRaFT 53.3 9.6 34 74.1 12.7 21.4
-Table 3: Experimental Results in MCQA Task.
 6.2 Ablation Experiments
@@
 We conduct sufficient ablation experiments to deeply investigate the impact of the above three factors on RAIT data selection.
-Due to space limitations, only the results of ablation experiments with the LLaMA2-7B-Chat model are reported, as shown in Table 4 and Table 5.
-Compared to Cor-RAIT, the method only introducing response certainty which named as Cor-RAIT w/o Flow achieved significant gains on the THS in the MCQA and OEQA tasks.
+Compared to Cor-RAIT, the method only introducing response certainty which named as “w/o Flow” achieved significant gains on the THS in the MCQA and OEQA tasks.
 This indicates that eliminating static conflicts can effectively mitigate the over-refusal of LLMs and this improvement is generalizable.
-Cor-RAIT w/o Cer only uses response correctness and knowledge flow.
+“w/o Cer” only uses response correctness and knowledge flow.
 Similarly, experimental results show that introducing knowledge flow to filter dynamic conflicts can also maintain the factuality of the model while improving its rejection capability.
 Finally, CRaFT considers both static and dynamic conflicts, further enhancing performance improvement.
-TriviaQA (ID) NQ (OOD)
-[…]
-Cor-RAIT 18.32 19.73 27.59 4.71
-CRaFT w/o Flow 20.45 21.12 38.75 5.15
-CRaFT w/o Cer 15.65 20.11 22.13 5.42
-CRaFT 12.93 23.29 22.38 6.54
-Table 4: Ablation Experiment Results in OEQA Task.
-MMLU (ID) ARC-c (OOD)
-[…]
-Cor-RAIT 39.57 11.32 38.14 11.08
-CRaFT w/o Flow 31.01 12.97 32.34 13.45
-CRaFT w/o Cer 32.05 11.51 32.94 13.85
-CRaFT 36.39 12.46 35.92 12.58
-Table 5: Ablation Experiment Results in MCQA Task
 7 Conclusion
-In this paper, we reveal that the over-refusal issue in existing correctness-based RAIT methods stems from the static and dynamic conflicts within the RAIT data.
-To address these, we introduced CRaFT: Firstly, by incorporating response certainty, we mitigate static conflicts during the construction of RAIT data.
-Secondly, through rehearsal training, we capture the trends of knowledge flow within the LLM, thereby overcoming dynamic conflicts.
-We conducted extensive experiments on the MCQA and OEQA tasks.
-The results show that CRaFT significantly outperforms the existing correctness-based baselines, thereby validating the effectiveness of our method.
-In the future, we plan to enhance CraFT by integrating RL-based training strategies.
-We also plan to explore its adaptation for more complex tasks, including reasoning and multi-turn dialogue.
-8 Acknowledgments
+In this paper, we identify over-refusal in correctness-based RAIT methods, caused by static and dynamic conflicts in RAIT data.
+To address this, we propose CRaFT: it mitigates static conflicts by incorporating response certainty during data construction and overcomes dynamic conflicts through rehearsal training to capture knowledge flow trends in LLMs.
+Extensive experiments on MCQA and OEQA tasks show CRaFT outperforms existing baselines, validating its effectiveness.
+Future work includes enhancing CRaFT with RL-based strategies and adapting it for more complex tasks, such as reasoning and multi-turn dialogue.
+Acknowledgments
 This research was supported by Shanghai Artificial Intelligence Laboratory.
@@
 (Xu et al.
-2024) defines precision (Prec) as $\frac{N_{c}}{N_{c}+N_{w}}$, where $N_{c}$, $N_{r}$, and $N_{w}$ represent the counts of correct, rejected, and wrong responses, respectively.
+2024a) defines precision (Prec) as $\frac{N_{c}}{N_{c}+N_{w}}$, where $N_{c}$, $N_{r}$, and $N_{w}$ represent the counts of correct, rejected, and wrong responses, respectively.
 Accuracy (Acc) is defined as $\frac{N_{c}}{N}$, and truthfulness (Truth) is given by $1-\frac{N_{w}}{N}$.